Find 36 ÷ 0.2 Find 36 ÷ 0.2 Find 36 ÷ 0.2 Find 36 ÷ 0.2

Find
 36 ÷ 0.2
Find
 36 ÷ 0.2
Find
 36 ÷ 0.2
Find
 36 ÷ 0.2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have,
= 36 ÷ (2/10)
= 36 × (10/2)
= 18 × 10
= 180

Thanks